# KYUSHU-HET-349

### v1

@@ 1 Introduction
Rare flavor-changing neutral current (FCNC) processes such as those involving $s\to d$ and $b\to s$ transitions are among the most sensitive probes of physics beyond the Standard Model (SM). Their theoretical cleanliness—stemming from the dominance of short-distance dynamics, and the absence of large hadronic uncertainties—makes them ideal for testing new physics scenarios.
Experimental efforts by NA62 [49] and KOTO [5] in Kaon decays, and Belle II [4] in $B$-meson decays, have significantly improved bounds on $K\to\pi\nu\bar{\nu}$ and $B\to K^{(*)}\nu\bar{\nu}$ channels, respectively. While the SM predicts a coherent pattern for these processes, emerging data provide mild yet intriguing hints of deviations. In particular, the central value of the $K^{+}\to\pi^{+}\nu\bar{\nu}$ branching ratio reported by NA62 lies somewhat above the SM prediction [49]. Even more strikingly, the latest Belle II results in the $B^{+}\to K^{+}\nu\bar{\nu}$ channel show an excess with a significance of about $2.7\sigma$ [4].
Many studies have investigated new physics interpretations of the recent Belle II excess in $B^{+}\to K^{+}\nu\bar{\nu}$ [4], mostly in two approaches. In the first approach, the excess is attributed to heavy new physics. Integrating out new heavy degrees of freedom induces various effective $bs\nu\nu$ interaction operators, which provide additional contributions to the $b\to s\nu\bar{\nu}$ transition. Effective field theories, with or without light right handed neutrinos, are employed to perform a model-independent analysis in this scenario [13, 17, 8, 59, 71, 43, 73, 85, 90, 35, 75, 9, 19, 78], allowing one to identify distinct correlations among various flavor observables. In particular, correlations among ${\cal B}(B^{+}\to K^{+}\nu\bar{\nu})$, ${\cal B}(B\to K^{*}\nu\bar{\nu})$, and longitudinal polarization fraction of $K^{*}$ in $B\to K^{*}\nu\bar{\nu}$ helps disentangle different new physics contributions to $B\to K^{(*)}\nu\bar{\nu}$ observables [17, 8, 73, 90, 35]. Concrete ultraviolet realizations are also proposed, including a supersymmetric model [92], models with leptoquarks [41, 85, 68, 93, 40, 52, 19] and $Z^{\prime}$ bosons [13, 85, 36], scotogenic model [42], grand unified theory [26], and universal extra-dimension model [91].
@@
MFV [44, 67, 33, 54] postulates that all sources of flavor and CP violation originate from the Standard Model Yukawa couplings, ensuring that any new physics sector obeys the same flavor structure as the SM. In this setup, new operators contributing to $s\to d\nu\bar{\nu}$ and $b\to s\nu\bar{\nu}$ transitions are aligned with CKM hierarchies, naturally generating correlated predictions for Kaon and $B$-meson observables. Moreover, MFV symmetries can stabilize additional flavor multiplets, allowing them to play the role of dark matter.
@@ 2 Dark Matter within Minimal Flavor Violation
The SM exhibits a large global flavor symmetry ${\rm U}(3)^{5}={\rm U}(3)_{q}\times{\rm U}(3)_{u}\times{\rm U}(3)_{d}\times{\rm U}(3)_{\ell}\times{\rm U}(3)_{e}$ in the gauge sector. This symmetry is explicitly broken by the Yukawa interactions of quarks and leptons to the doublet Higgs field $H$, and flavor violating processes show specific patterns controlled by the structure of the Yukawa interaction matrices $Y_{u,d,e}$.
@@
In this framework, interactions between quarks and dark matter particles determine the phenomenology in the dark sector: e.g. dark matter production in the early universe, for (in)direct searches, and decays of the heavy dark states. The traditional WIMP parameter space, where $\chi$ is heavier than $B$ mesons, was comprehensively studied in [15, 83], with implications ranging from freeze-out cosmology to collider signals and FCNC processes. The possibility of multi-component dark matter was investigated in [87]. If instead $\chi$ is lighter than $B$ mesons, some operators can induce dark matter emission from heavy quark decays, such as $d_{i}\to d_{j}\chi_{i}\bar{\chi}_{j}$. Since $\chi$ must be effectively unobservable with current detectors, these new decay modes mimic the FCNC decays, $d_{i}\to d_{j}\nu\bar{\nu}$.
@@ Acknowledgments
We thank Ludovico Vittorio for useful discussions at the early stage of this work. The work of C.T. has received funding from the French ANR, under contracts ANR-19-CE31-0016 (‘GammaRare’) and ANR-23-CE31-0018 (‘InvISYble’), that he gratefully acknowledges. The work of S.O. is supported by JSPS KAKENHI Grant No. JP25K17401 and by an appointment to the JRG Program at the APCTP through the Science and Technology Promotion Fund and Lottery Fund of the Korean Government and by the Korean Local Governments – Gyeongsangbuk-do Province and Pohang City.
@@ A.1 Belle II
This expression applies to either the ITA or HTA dataset, where $n^{i}_{\text{DM}}$, $n^{i}_{\nu\bar{\nu}}$, and $n^{i}_{\rm bkg}$ are the yields in the $i^{\rm th}$ bin for the signal, the $B^{+}\to K^{+}\nu\bar{\nu}$ decay, and the background, respectively. Within the two methods, the first two components on the r.h.s. of Eq. (A.4) are defined as
@@ A.3 Belle
In the case of Belle search, we have no access to the kinematic distribution of the events and, additionally, their measurements only put upper limits on the decay rates. Therefore we construct a likelihood term that constraints only the total BR of the DM signal such as
${\mathcal{L}}_{\text{BaBar,k}}^{(2)}\propto\prod_{k}\text{exp}\left\{-\frac{1}{2}\left(\frac{{\cal B}_{\text{DM}}}{\sigma_{{\cal B}_{\text{DM}}}}\right)^{2}\right\}$ (A.14)
@@
with ${\cal B}_{\text{exp}}$ the experimental limit in Eqs. (3.6)-(3.9) and $\chi^{2}(90\%,1\text{dof})$ the value of a chi-squared variable at 90%CL with 1 degree of freedom.
@@ A.5 KOTO
The likelihood for the KOTO measurement is constructed similarly to the one for NA62. In this case, acpording to Ref. [5], we set $n_{\text{obs}}=0$, $\mu_{b}=0.252$ and $\sigma_{b}=0.08$, where the latter is obtained by averaging over the asymmetric statistical uncertainty and then adding in quadrature the statistical and systematic uncertainties.
@@ A.6 SM prediction of neutrino modes
The general theoretical expressions of the rare meson decay rates with neutrino or DM emission can be found in Appendix B. These expressions depend on form factors which are introduced to parametrize the hadronic matrix elements of operators involving quark fields, see formulae in Appendix C. We will take the needed form factors from the most recent extrapolation based on lattice or experimental data available in literature, see Appendix C for details. In this work we will neglect the uncertainties related the form factors of the decay rates.
@@
However we will consider the uncertainty on the SM prediction of the total BR for the $B\to K^{(*)}\nu\bar{\nu}$ and $K\to\pi\nu\bar{\nu}$’s BRs, whose values are taken recollected in Table 4. We hence define a nuisance parameter for each of them as

### v2

@@ 1 Introduction
Rare flavor-changing neutral current (FCNC) processes such as those involving $s\to d$ and $b\to s$ transitions are among the most sensitive probes of physics beyond the Standard Model (BSM). Their theoretical cleanliness—stemming from the dominance of short-distance dynamics, and the absence of large hadronic uncertainties—makes them ideal for testing new physics scenarios.
Experimental efforts by NA62 [49] and KOTO [5] in Kaon decays, and Belle II [4] in $B$-meson decays, have significantly improved bounds on $K\to\pi\nu\bar{\nu}$ and $B\to K^{(*)}\nu\bar{\nu}$ channels, respectively. While the Standard Model (SM) predicts a coherent pattern for these processes, emerging data provide mild yet intriguing hints of deviations. In particular, the central value of the $K^{+}\to\pi^{+}\nu\bar{\nu}$ branching ratio reported by NA62 lies somewhat above the SM prediction [49]. Even more strikingly, the latest Belle II results in the $B^{+}\to K^{+}\nu\bar{\nu}$ channel show an excess with a significance of about $2.7\sigma$ [4].
Many studies have investigated new physics interpretations of the recent Belle II excess in $B^{+}\to K^{+}\nu\bar{\nu}$ [4], mostly in two approaches. In the first approach, the excess is attributed to heavy new physics. Integrating out new heavy degrees of freedom induces various effective $bs\nu\nu$ interaction operators, which provide additional contributions to the $b\to s\nu\bar{\nu}$ transition. Effective field theories (EFTs), with or without light right handed neutrinos, are employed to perform a model-independent analysis in this scenario [13, 17, 8, 59, 72, 43, 74, 86, 91, 35, 76, 9, 19, 79], allowing one to identify distinct correlations among various flavor observables. In particular, correlations among ${\cal B}(B^{+}\to K^{+}\nu\bar{\nu})$, ${\cal B}(B\to K^{*}\nu\bar{\nu})$, and longitudinal polarization fraction of $K^{*}$ in $B\to K^{*}\nu\bar{\nu}$ helps disentangle different new physics contributions to $B\to K^{(*)}\nu\bar{\nu}$ observables [17, 8, 74, 91, 35]. Concrete ultraviolet realizations are also proposed, including a supersymmetric model [93], models with leptoquarks [41, 86, 69, 94, 40, 52, 19] and $Z^{\prime}$ bosons [13, 86, 36], scotogenic model [42], grand unified theory [26], and universal extra-dimension model [92].
@@
MFV [44, 68, 33, 54] postulates that all sources of flavor and CP violation originate from the Standard Model Yukawa couplings, ensuring that any new physics sector obeys the same flavor structure as the SM. In this setup, new operators contributing to $s\to d\nu\bar{\nu}$ and $b\to s\nu\bar{\nu}$ transitions are aligned with Cabibbo–Kobayashi–Maskawa (CKM) hierarchies, naturally generating correlated predictions for Kaon and $B$-meson observables. Strikingly, MFV symmetries can stabilize additional flavor multiplets, allowing them to play the role of dark matter. It has previously been shown that in MFV the lightest state of a new QCD-singlet field $\chi$ that transforms under the quark flavor subgroup ${\rm SU}(3)_{q}\times{\rm SU}(3)_{u}\times{\rm SU}(3)_{d}$ is stable, even when all higher dimensional operators are included [88, 15]. This is due only to the invariance under the color and flavor groups within the MFV. Remarkably, the addition of the $\chi$ fields, acting as spurions of the flavor group beyond the usual SM Yukawa matrices, breaks the minimally expected scaling of BSM contribution in strict U(3)3 MFV: $C^{\text{BSM}}_{ij}\sim(Y_{u}^{\dagger}Y_{u})_{ij}\propto V^{*}_{ti}V_{tj}$. Investigation of MFV beyond minimal assumptions is therefore well motivated, in particular considering implications for $d_{i}\rightarrow d_{j}\chi_{i}\bar{\chi}_{j}$ processes, where $\chi_{i}$ has the same flavor charge as the quark fields.
@@ 2 Dark Matter within Minimal Flavor Violation
The SM exhibits a large global flavor symmetry ${\rm U}(3)^{5}={\rm U}(3)_{q}\times{\rm U}(3)_{u}\times{\rm U}(3)_{d}\times{\rm U}(3)_{\ell}\times{\rm U}(3)_{e}$ in the gauge sector [44, 63]. This symmetry is explicitly broken by the Yukawa interactions of quarks and leptons to the doublet Higgs field $H$, and flavor violating processes show specific patterns controlled by the structure of the Yukawa interaction matrices $Y_{u,d,e}$.
@@
In this framework, interactions between quarks and dark matter particles determine the phenomenology in the dark sector: e.g. dark matter production in the early universe, for (in)direct searches, and decays of the heavy dark states. The traditional Weakly Interacting Massive Particle (WIMP) parameter space, where $\chi$ is heavier than $B$ mesons, was comprehensively studied in [15, 84], with implications ranging from freeze-out cosmology to collider signals and FCNC processes. The possibility of multi-component dark matter was investigated in [88]. If instead $\chi$ is lighter than $B$ mesons, some operators can induce dark matter emission from heavy quark decays, such as $d_{i}\to d_{j}\chi_{i}\bar{\chi}_{j}$. Since $\chi$ must be effectively unobservable with current detectors, these new decay modes mimic the FCNC decays, $d_{i}\to d_{j}\nu\bar{\nu}$.
@@ Acknowledgments
We thank Ludovico Vittorio for useful discussions at the early stage of this work and Jernej Kamenik for providing further insights on the analysis of Ref. [29]. The work of C.T. has received funding from the French ANR, under contracts ANR-19-CE31-0016 (‘GammaRare’) and ANR-23-CE31-0018 (‘InvISYble’), that he gratefully acknowledges. The work of S.O. is supported by JSPS KAKENHI Grant No. JP25K17401 and by an appointment to the JRG Program at the APCTP through the Science and Technology Promotion Fund and Lottery Fund of the Korean Government and by the Korean Local Governments – Gyeongsangbuk-do Province and Pohang City. The work of FM is supported by the European Union-Next Generation EU and by the Italian Ministry of University and Research (MUR) via the PRIN 2022 project No. 2022K4B58X-AxionOrigins. This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology). We also wish to thank the Theory Group of the Laboratori Nazionali di Frascati for their hospitality during the first stage of this work.
@@ A.1 Belle II
This expression applies to either the ITA or HTA dataset, where $n^{i}_{\text{DM}}$, $n^{i}_{\nu\bar{\nu}}$, and $n^{i}_{\rm bkg}$ are the yields in the $i^{\rm th}$ bin for the signal, the $B^{+}\to K^{+}\nu\bar{\nu}$ decay, and the background, respectively. Within the two methods, the first two components on the right hand side of Eq. (A.4) are defined as
@@ A.3 Belle
In the case of Belle search, we have no access to the kinematic distribution of the events and, additionally, their measurements only put upper limits on the decay rates. Therefore we construct a likelihood term that constrains only the total branching ratio (BR) of the DM signal such as
${\mathcal{L}}_{\text{Belle}}^{(2)}\propto\prod_{k}\text{exp}\left\{-\frac{1}{2}\left(\frac{{\cal B}_{\text{DM}}}{\sigma_{{\cal B}_{\text{DM}}}}\right)^{2}\right\}$ (A.14)
@@
with ${\cal B}_{\text{exp}}$ the experimental limit in Eqs. (3.6)-(3.9), ${\cal B}_{\nu\bar{\nu}}$ the SM prediction of the di-neutrino channel and $\chi^{2}(90\%,1\text{dof})$ the value of a chi-squared variable at 90%CL with 1 degree of freedom. The numerator quantifies the constraint on possible additional BSM contributions, while the denominator accounts for rescaling to a $68\%$ confidence interval, such that Eq. (A.14) is a Gaussian constraint term.
@@ A.5 KOTO
The likelihood for the KOTO measurement is constructed similarly to the one for NA62. In this case, according to Ref. [5], we set $n_{\text{obs}}=0$, $\mu_{b}=0.252$ and $\sigma_{b}=0.08$, where the latter is obtained by averaging over the asymmetric statistical uncertainty and then adding in quadrature the statistical and systematic uncertainties.
@@ A.6 SM prediction of neutrino modes
The general theoretical expressions of the rare meson decay rates with neutrino or DM emission can be found in Appendix B. These expressions depend on form factors which are introduced to parametrize the hadronic matrix elements of operators involving quark fields, see formulae in Appendix C. We will take the needed form factors from the most recent extrapolation based on lattice or experimental data available in literature, see Appendix C for details. In this work we will neglect the uncertainties related to the form factors of the decay rates, in practical terms this means that uncertainties on the expected shape of the $q^{2}$ distribution are neglected.
@@
In contrast, we introduce a likelihood term to take into account the uncertainties on the SM prediction of the total BRs for the $B\to K^{(*)}\nu\bar{\nu}$ and $K\to\pi\nu\bar{\nu}$, which are taken from literature as reported in Table 4. We hence define a nuisance parameter for each of them as
@@
On a practical level, we found that the uncertainty on the SM predictions of the neutrino modes has negligible impact on our numerical results.